# Towards Building the Federated GPT: Federated Instruction Tuning

### v1

@@ Abstract
While “instruction-tuned" generative large language models (LLMs) have demonstrated an impressive ability to generalize to new tasks, the training phases heavily rely on large amounts of diverse and high-quality instruction data (such as ChatGPT and GPT-4). Unfortunately, acquiring high-quality data, especially when it comes to human-written data, can pose significant challenges both in terms of cost and accessibility. Moreover, concerns related to privacy can further limit access to such data, making the process of obtaining it a complex and nuanced undertaking. Consequently, this hinders the generality of the tuned models and may restrict their effectiveness in certain contexts. To tackle this issue, our study introduces a new approach called Federated Instruction Tuning (FedIT), which leverages federated learning (FL) as the learning framework for the instruction tuning of LLMs. This marks the first exploration of FL-based instruction tuning for LLMs. This is especially important since text data is predominantly generated by end users. For example, collecting extensive amounts of everyday user conversations in different languages can be a useful approach to improving the generalizability of LLMs, allowing them to generate authentic and natural responses. Therefore, it is imperative to design and adapt FL approaches to effectively leverage these users’ diverse instructions stored on local devices, while preserving privacy and ensuring data security. In the current paper, by conducting widely used GPT-4 auto-evaluation, we demonstrate that by exploiting the heterogeneous and diverse sets of instructions on the client’s end with the proposed framework FedIT, we improved the performance of LLMs compared to centralized training with only limited local instructions. Further, in this paper, we developed a Github repository named Shepherd. This repository offers a foundational framework for exploring federated fine-tuning of LLMs using heterogeneous instructions across diverse categories. The framework is designed for ease of use, adaptability, and scalability to accommodate large datasets. Additionally, it facilitates the seamless integration of novel algorithms and configurations, making it a convenient tool for researchers and practitioners in the NLP community.
@@ 2.2 Federated Learning in NLP Tasks
Federated Learning [47] is a decentralized and collaborative machine learning technique that enables data to remain on user devices. Significant research efforts have focused on addressing privacy and heterogeneity challenges and developing advanced FL methods [26, 44, 51, 80]. These advancements include designing optimization methods with improved aggregation performance ([9, 58, 60, 70, 85], increasing the framework’s robustness against adversarial attacks [64], devising effective client selection mechanisms [11, 19, 69, 82], enhancing personalization capabilities [14, 33, 71], and boosting the overall efficiency of FL systems [30, 32, 46, 59].
@@ 4.1 Implementation details
In our FL setup, we assume the presence of 100 clients. We proceed to apply the Shepherd framework’s second data partitioning technique to divide the residual data from the Databricks-dolly-15k dataset into 100 distinct portions. Each of these portions corresponds to an individual client’s local instruction dataset. We conduct a total of 20 communication rounds, with each round involving the random selection of 5 (0.05%) clients for training. Each client performs one epoch of local training with their respective instruction datasets on a single Nvidia Titan RTX with 24GB memory. We initialize the model with the 7B LLaMA model. The model remains frozen during training, thereby reducing GPU memory usage and enhancing training speed. In alignment with Baize’s settings [77], we apply LoRA to all linear layers with a rank of 8 to boost adaptation capabilities. Following [24], we use random Gaussian initialization for A and set B to zero, ensuring that the value of BA is zero at the beginning of training. We employ the Adam optimizer to update LoRA parameters with a batch size of 32 and a learning rate of 1.5e-4. We set the maximum input sequence length to 512 and provide the template of the prompt adopted from Alpaca-lora in Table 1. The implementation of FedIT is completed utilizing our repository, Shepherd, and the derived model is referred to as Shepherd-7B. We detail the number of model parameters, training time, and GPU memory consumption in Table 2.
@@ 4.2 Qualitative Study with Automatic Evaluation
As demonstrated in Table 3, the performance of our proposed model, Shepherd-7B, significantly surpasses that of the LLaMA model. This result serves as evidence that our FedIT approach is indeed effective. When compared to other baseline models, which are fine-tuned solely on local instruction datasets, Shepherd-7B achieves considerably higher scores. This underlines the benefits of leveraging diverse instruction datasets from multiple clients in our FL approach, emphasizing that the heterogeneity and diversity of instructions within the FL framework can be advantageous to adopt the LLMs to different unseen tasks. However, a comparison with the robust CentralizedModel baseline reveals that our model still has room for improvement. This disparity is partly attributed to the fact that the local models aggregated at the server side are trained on instructions with substantially different distributions, which can cause their local models to learn different representations of the instructions. Consequently, there is a need for further exploration of more efficient federated optimization methods to enhance the aggregation process. In conclusion, as discussed in Section 3.2, statistical heterogeneity can be a beneficial factor for FedIT, as it enhances the diversity of instruction data, thus improving the model’s generalization ability to unseen tasks. However, to fully utilize the benefits of data heterogeneity, advanced federated optimization methods need to be developed and integrated to manage and leverage heterogeneity more effectively.
@@ 5.3 Personalization
With deploying FL in LLM, due to the differences among the language data (instructions) used in distributed clients and averaging of learning updates across a decentralized population, personalization becomes a critical requirement for FL systems [43]. The former can be further complicated by language diversity, domain-specific instructions, task complexity, emotional tone, cultural factors, etc., which are new aspects of heterogeneity [39, 76]. For instance, in multilingual applications, fairness across languages, especially for languages with fewer data samples, is essential but hard to achieve[73, 76]. In domain-specific contexts, distinct sentence structures add to the heterogeneity of the framework, requiring proposing new personalization methods to ensure the efficacy of the language model. Methods that combine personal embeddings with shared context embeddings, and preference embeddings, that facilitate personalization without the need for backpropagation, etc. have the potential to revolutionize the field of NLP.

### v2

@@ Abstract
While "instruction-tuned" generative large language models (LLMs) have demonstrated an impressive ability to generalize to new tasks, the training phases heavily rely on large amounts of diverse and high-quality instruction data (such as ChatGPT and GPT-4). Unfortunately, acquiring high-quality instructions, especially when it comes to human-written instructions, can pose significant challenges both in terms of cost and accessibility. Moreover, concerns related to privacy can further limit access to such data, making the process of obtaining it a complex and nuanced undertaking. Consequently, this hinders the generality of the tuned models and may restrict their effectiveness in certain contexts. To tackle this issue, our study introduces a new approach called Federated Instruction Tuning (FedIT), which leverages federated learning (FL) as the learning framework for the instruction tuning of LLMs. This marks the first exploration of FL-based instruction tuning for LLMs. This is especially important since text data is predominantly generated by end users. For example, collecting extensive amounts of everyday user conversations in different languages can be a useful approach to improving the generalizability of LLMs, allowing them to generate authentic and natural responses. Therefore, it is imperative to design and adapt FL approaches to effectively leverage these users’ diverse instructions stored on local devices, while preserving privacy and ensuring data security. In the current paper, by conducting widely used GPT-4 auto-evaluation, we demonstrate that by exploiting the heterogeneous and diverse sets of instructions on the client’s end with the proposed framework FedIT, we improved the performance of LLMs compared to centralized training with only limited local instructions. Further, in this paper, we developed a Github repository named Shepherd. This repository offers a foundational framework for exploring federated fine-tuning of LLMs using heterogeneous instructions across diverse categories. The framework is designed for ease of use, adaptability, and scalability to accommodate large datasets. Additionally, it facilitates the seamless integration of novel algorithms and configurations, making it a convenient tool for researchers and practitioners in the NLP community.
@@ 2.2 Federated Learning in NLP Tasks
Federated Learning [55] is a decentralized and collaborative machine learning technique that enables data to remain on user devices. Significant research efforts have focused on addressing privacy and heterogeneity challenges and developing advanced FL methods [33, 52, 59, 92, 93]. These advancements include designing optimization methods with improved aggregation performance ([10, 21, 28, 66, 68, 79, 102], increasing the framework’s robustness against adversarial attacks [72], devising effective client selection mechanisms [12, 25, 78, 95], enhancing personalization capabilities [18, 40, 80, 89], and boosting the overall efficiency of FL systems [37, 39, 54, 67, 74].
@@ 4.1 Implementation details
In our FL setup, we assume the presence of 100 clients. We proceed to apply the Shepherd framework’s second data partitioning technique to divide the residual data from the Databricks-dolly-15k dataset into 100 distinct portions. Each of these portions corresponds to an individual client’s local instruction dataset. We conduct a total of 20 communication rounds, with each round involving the random selection of 5 (5%) clients for training. Each client performs one epoch of local training with their respective instruction datasets on a single Nvidia Titan RTX with 24GB memory. We initialize the model with the 7B LLaMA model. The model remains frozen during training, thereby reducing GPU memory usage and enhancing training speed. In alignment with Baize’s settings [88], we apply LoRA to all linear layers with a rank of 8 to boost adaptation capabilities. Following [31], we use random Gaussian initialization for A and set B to zero, ensuring that the value of BA is zero at the beginning of training. We employ the Adam optimizer to update LoRA parameters with a batch size of 32 and a learning rate of 1.5e-4. We set the maximum input sequence length to 512 and provide the template of the prompt adopted from Alpaca-lora in Table 1. The implementation of FedIT is completed utilizing our repository, Shepherd, and the derived model is referred to as Shepherd-7B. We detail the number of model parameters, training time, and GPU memory consumption in Table 2.
@@ 4.2 Qualitative Study with Automatic Evaluation
As demonstrated in Table 3, the performance of our proposed model, Shepherd-7B, significantly surpasses that of the LLaMA model. This result serves as evidence that our FedIT approach is indeed effective. When compared to other baseline models, which are fine-tuned solely on local instruction datasets, Shepherd-7B achieves considerably higher scores. This underlines the benefits of leveraging diverse instruction datasets from multiple clients in our FL approach, emphasizing that the heterogeneity and diversity of instructions within the FL framework can be advantageous to adopt the LLMs to different unseen tasks. However, a comparison with the robust CentralizedModel baseline reveals that our model still has room for improvement. This disparity is partly attributed to the fact that the local models aggregated at the server side are trained on instructions with substantially different distributions, which can cause their local models to learn different representations of the instructions.Consequently, there is a need for further exploration of more efficient federated optimization methods and client scheduling methods, such as FA-LD [16] and FedCBS [95], which are inspired by bayesian sampling methods [47, 87, 96, 97, 99, 100], to enhance the aggregation process. In conclusion, as discussed in Section 3.2, statistical heterogeneity can be a beneficial factor for FedIT, as it enhances the diversity of instruction data, thus improving the model’s generalization ability to unseen tasks. However, to fully utilize the benefits of data heterogeneity, advanced federated optimization methods need to be developed and integrated to manage and leverage heterogeneity more effectively.
@@ 5.3 Personalization
With deploying FL in LLM, due to the differences among the language data (instructions) used in distributed clients and averaging of learning updates across a decentralized population, personalization becomes a critical requirement for FL systems [51]. The former can be further complicated by language diversity, domain-specific instructions, task complexity, emotional tone, cultural factors, etc., which are new aspects of heterogeneity [46, 86]. For instance, in multilingual applications, fairness across languages, especially for languages with fewer data samples, is essential but hard to achieve [82, 86]. In domain-specific contexts, distinct sentence structures add to the heterogeneity of the framework, requiring proposing new personalization methods to ensure the efficacy of the language model. Methods, such as meta-learning [2, 14, 22] or few-shot learning [15, 83], that combine personal embeddings with shared context embeddings, and preference embeddings, that facilitate personalization without the need for backpropagation, etc. have the potential to revolutionize the field of NLP.